OpenStudy (anonymous):

Solve for the variable: p – 4.8 6

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

if you mean p-4.8=6, then p=6+4.8 p=10.8